# SourceNet: Interpretable Sim-to-Real Inference on Variable-Geometry Sensor Arrays for Earthquake Source Inversion

### v1

@@ Abstract
Inferring high-dimensional physical states from sparse, ad-hoc sensor arrays is a fundamental challenge across AI for Science, as they are complicated by irregular geometries and the profound Sim-to-Real gap in physical modeling. Taking earthquake source characterization as a representative challenge, we address limitations in conventional deep learning: CNNs demand fixed grids, while pooling-based architectures (e.g., DeepSets) struggle to capture the relational wave physics. Here, we propose SourceNet, a Transformer-based framework that treats the sensor array as a flexible set to model arbitrary geometries. To bridge the reality gap, we introduce Physics-Structured Domain Randomization (PSDR). Instead of forcing feature alignment, PSDR randomizes the governing physical dynamics by varying velocity structures, propagation effects, and sensor availability, to force the model to learn robust representations invariant to unmodeled environmental heterogeneity. By pre-training on 100,000 synthetic events and fine-tuning on $\sim$2,000 real world events, SourceNet achieves state-of-the-art precision on held-out real data. This demonstrates exceptional data efficiency, and matches classical solvers while enabling real-time processing. Remarkably, interpretability analysis reveals that the model shows scientific-agent-like features: it autonomously discovers geometric information bottlenecks and learns an attention policy that prioritizes sparse sensor placements, effectively recovering principles of optimal experimental design from data alone.
@@ 1 Introduction
How do we reconstruct a complex physical event, whether it is a distant earthquake, a cosmic radio burst, or a climate anomaly, using only a handful of scattered sensors? This is the archetypal challenge across the physical sciences (Bergen et al., 2019; Bouman et al., 2016; Ravuri et al., 2021): inferring complex physical source parameters $\theta$ from a sparse set of indirect, noisy observations $X=\{x_{1},\dots,x_{n}\}$. Unlike images or video where data pixels are fixed on a regular grid, scientific data is often collected by ad-hoc sensor arrays where the number of sensors varies per event and their spatial arrangement is irregular.
@@
Compounding the geometric challenge is the Sim-to-Real gap (Tobin et al., 2017). In science disciplines like geophysics and robotics, obtaining ground-truth labels for real-world observations is expensive and labor-intensive, therefore models have to trained on physics-based simulations, which inevitably lack the dirty reality including instrumental noise, realistic wave propagation effect, and structural heterogeneity (Karniadakis et al., 2021). Hence, naive training on idealized simulations faces a substantial distribution shift and is difficult to generalize to various real-world data. A trustworthy AI for Science framework must learn invariant physical operations that hold true despite these distribution shifts.
@@
    We bridge the Sim-to-Real gap not by feature alignment, but by randomizing the physics itself (e.g., velocity models, sensor dropouts), forcing the model to learn representations invariant to crustal heterogeneity.
@@ 2.1 Deep Learning on Sets and Irregular Grids
Handling varying cardinality inputs require permutation-invariant architectures. Early set-based architectures, such as DeepSets (Zaheer et al., 2017) and PointNet (Qi et al., 2017), can achieve permutation invariance by applying independent encoders followed by symmetric global pooling. While effective for simple classification tasks, we argue these are insufficient for wave physics because pooling reduces the set to a sum, which limits their ability to model relational context (e.g., amplitude differences between stations) critical for inversion. Graph Neural Networks (GNNs) (Kipf, 2016; Battaglia et al., 2018) model dependencies explicitly but typically assume fixed edges or require expensive dynamic graph construction ($O(N^{2})$) (Wang et al., 2019). More recently, Neural Operators (e.g., DeepONet(Lu et al., 2021), FNO(Li et al., 2020)) have emerged to learn mappings between infinite-dimensional function spaces. However, these methods often assume dense sampling of fixed control points. In contrast, Set Transformers (Lee et al., 2019) leverage self-attention to model pairwise interactions dynamically. In the context of wave physics, attention-based mechanisms offer a theoretical advantage over pooling-based baselines by preserving radiation patterns on sparse arrays.
@@ 2.2 Sim-to-Real Transfer and Domain Randomization
Bridging the simulation-to-reality gap is a key challenge of domains ranging from robotics (Tobin et al., 2017) and neurobioscience (Gonçalves et al., 2020) to geophysics. In these fields, simulations provide abundant data but inevitably simplify the complex, unmodeled dynamics of the physical world (Pan & Yang, 2009). Traditional unsupervised domain adaptation methods aim to align feature distributions via adversarial training (Ganin et al., 2016) or minimizing maximum mean discrepancy (Long et al., 2015), but these require direct access to the target domain during training and can be unstable. Domain randomization (DR) (Tobin et al., 2017) takes a different approach: it randomizes simulator parameters to force the model to learn invariants (Peng et al., 2018). While DR has become standard in robotics for handling visual and contact dynamics (Akkaya et al., 2019), scientific inverse problems require a deeper form of adaptation because the uncertainty lies in the physics themselves. Our SourceNet extends DR by randomizing the governing physical processes (velocity structure, wave propagation effect, scattering) rather than just visual appearance. This not only solves the specific challenge of seismic inversion but offers a generalizable blueprint for learning robust operators in many physical system with imperfect simulations.
@@ 3 Problem Formulation and Methodology
The physical source ($y$). Our target is the seismic event $\mathcal{E}$ parameterized by a latent physical state $\mathbf{y}\in\mathbb{R}^{6}$. This state encapsulates the Moment Tensor (3*3 symmetric tensor) (Aki & Richards, 2002), representing the fault orientation and slip direction (5 independent zero-trace $M_{ij}$ components), and the moment magnitude $M_{w}$ indicating the energy scale (Kanamori, 1977).
The Observation Set ($X$). Unlike images defined on regular grids, a seismic network is an ad-hoc collection of sensors. We represent the observation as an unordered set $\mathbf{X}=\{(\mathbf{w}_{i},\mathbf{s}_{i})\}_{i=1}^{N}$, where $N$ is the number of stations (varies per event). Each element in the set is a multi-modal tuple: $\mathbf{w}_{i}\in\mathbb{R}^{C\times T}$ represents the dual-domain waveform features, including both the raw velocity time-series (for phase information) and their spectral representations (for frequency-dependent properties); $\mathbf{s}_{i}\in\mathbb{R}^{d}$ represents scalar metadata including station azimuth, epicentral distance, locations, and the amplitude and amplitude ratios of the waveforms. Thus, they encode the physical attenuation from distance-based amplitude decay necessary to unravel the source magnitude.
@@ 3.1 Physics-Structured Domain Randomization (PSDR)
  • Network Availability ($\mathcal{M}$): Real sensors can fail or have gaps. We apply a stochastic masking operator $\mathcal{M}$ (Bernoulli dropout) to randomly discard stations during training, to make the model robust to sparse and variable geometries.
Validation via Manifold Alignment. To validate our physics-structured approach, we compare the latent spaces of the baseline model trained on clean synthetics and SourceNet trained on PSDR, where both models are fine-tuned on the same subset of real data. As shown in the t-SNE (Maaten & Hinton, 2008) visualizations in Figure 3, standard fine-tuning is insufficient to bridge the reality gap, shown by the disjoint clusters of the naive baseline model, and it memorize real samples rather than learning generalized physics. In contrast, SourceNet achieves perfect manifold alignment. This proves that PSDR is a prerequisite for effective transfer learning, allowing the model to learn domain-invariant physical features that fine-tuning alone cannot recover.
@@ 3.2 SourceNet Architecture
SourceNet is designed to mirror how a seismologist would analyzes an event: first extracting features from each station (phases, amplitudes, location), and then synthesizing this information globally to constrain the source. As illustrated in Figure 1, the model consists of three functional modules:
@@
  • P-Wave Tower: We extract a 6-second window ([1s,+5s][-1s,+5s] relative to the P wave onset) filtered in the low-frequency band (0.1–2 Hz) to remove the drift and noise. This stream processes both the raw velocity waveforms and their spectral representations via a 1D ResNet (He et al., 2016). This stream captures not only the high-frequency onset polarity, but also the early crustal phases in the coda. As our analysis reveals, these converted phases contain rich constraints often ignored by traditional first-motion solvers.
@@
  • Scalar Tower: Standard Transformers (Vaswani et al., 2017) use discrete positional encodings (like integers in a sentence), which can fail for irregular sensor arrays. We instead project a vector of 20 explicit physical features through an MLP. This vector includes the geometric metadata: Station coordinates (lat/lon), azimuth, epicentral distance, and source depth; the amplitude information: Max absolute amplitudes for both time-domain and spectral channels (6 for P-wave, 6 for S-wave); and 3 P/S amplitude ratios. This generates a continuous embedding that allows the scale-invariant Transformer to recover the absolute Moment Magnitude ($M_{w}$) and the source radiation patterns.
@@
2. Global Context via Self-Attention. A challenge to the source inversion is that local observations are ambiguous: a single sensor cannot distinguish between different physical parameters without reference to other sensor data. Standard pooling (DeepSets) can fail because it aggregates these ambiguous features directly and masks the relational context. We employ a Transformer Encoder to model these relations. Through self-attention, each station queries the entire array, aggregating information from distant sensors to update its own representation. This process contextualizes the local data features against the global radiation pattern, which allows each station’s embedding to encode not just what it saw, but how it fits into the collective source mechanism.
@@ 4.1 Datasets
Real-world Fine-tuning Domain. We use a high-quality catalog of 2,544 $M>3.0$ earthquakes in Southern California with focal mechanism solutions from (Yang et al., 2012). These events are selected for their reliability (Quality A & B) and acceptable station coverage (¿5 stations) (Hardebeck & Shearer, 2002).
@@ 5 Experimental Results and Analysis
We evaluate SourceNet on both the held-out synthetic test set (for theoretical upper bound) and the real-world Southern California catalog (to measure Sim-to-Real transfer).
@@ 5.1 Sim-to-Real Generalization
As shown in Figure 4 (Top), SourceNet predicts source parameters very well on the synthetic test set, with Mean Absolute Errors (MAE) for moment tensor ($M_{ij}$) ranging from $0.05$ to $0.08$. This shows that the architecture can well approximate the nonlinear wave physics. When applied to real-world data (Figure 4, Bottom), we observe a degradation due to the inevitable reality gap. However, the predictions remain nicely correlated with ground truth, and the magnitude prediction remains robust ($MAE=0.11$) although real-world waveforms are distorted by scattering. As we demonstrate in Section 5.3, this robustness is directly attributable to our multi-modal design, which disentangles geometric spreading (via the scalar tower) from source physics.
Figure 4: Sim-to-Real Inference Performance. Evaluation on the held-out real-world catalog. (a) Predicted vs. true Magnitude ($M_{w}$) demonstrates high accuracy ($MAE=0.11$), confirming that the scalar tower successfully disentangles source energy from path attenuation. (b-c) Normalized Moment Tensor components ($M_{xx},M_{xy}$) maintain robust linearity despite the domain gap. (d) The Kagan angle error distribution (Mean $26.18^{\circ}$, Median $19.37^{\circ}$) confirms that SourceNet achieves SOTA precision and effectively hits the label uncertainty floor of the manual catalogs.
@@ 5.2 Benchmarking against SOTA
We benchmark SourceNet against established baselines: the authoritative Southern California catalog (Yang et al., 2012) and recent deep learning solvers. On our operational test set that mixes both good (quality A) and acceptable (quality B) events, SourceNet achieves a mean Kagan angle of $26.18^{\circ}$ and a median of $19.37^{\circ}$. Given that ground-truth uncertainties for Quality B events range up to $35^{\circ}$ (Yang et al., 2012), our error effectively hits the label noise floor. This suggests the model is not merely fitting noisy labels but has learned a robust physical operator more consistent than the noisy manual labels used for supervision.
@@ 5.3 Ablation Studies
To validate our architecture, we perform component-wise ablation studies on the synthetic test set. We train two variants: one removing the Scalar Tower (waveform only), and one replacing the Set Transformer with a DeepSets-style pooling architecture. The results are summarized in Table 1. We found that removing the scalar metadata causes the Magnitude prediction to collapse as MAE increasing by over 150% (0.07 $\to$ 0.18). This shows that waveform shapes alone are ambiguous on energy scale, and explicit geometric context (distance, attenuation) fed by the Scalar Tower is essential for disentangling energy from propagation effects. Also, we found replacing the Transformer with simple pooling (DeepSets) makes the mechanism estimation worse: the $M_{ij}$ error nearly triples (0.06 $\to$ 0.17), and the Kagan angle degrades significantly ($10.2^{\circ}\to 28.4^{\circ}$). This suggests that wave physics is relational: determining a source mechanism requires comparing relative data information across the array. Self-Attention captures these pairwise dependencies, whereas global pooling masks them.
@@ 6 Interpretability: Emergent Physical Strategies
Deep learning models are often criticized as black boxes (Rudin, 2019). Here, we show that SourceNet is not just fitting function, but has autonomously discovered physical strategies that extend human intuitions (Iten et al., 2020).
A question of interest is how much of a waveform is actually useful. Classical solvers typically rely on a single data point: the first P-wave polarity. The rest of the time series is often discarded as coda noise due to complex scattering. However, our Grad-CAM analysis (Figure 5) shows that SourceNet’s attention is not confined to the P-wave onset but extends substantially into the coda and early scattered phases. Physically, this suggests the model has learned that these phases are not random but contains rich information about the source mechanism (Aki & Richards, 2002). Thus, SourceNet effectively performs an implicit full-waveform inversion from the entire signal duration that human analysts usually ignore (Fichtner, 2010).
@@
This behavior represents an autonomous discovery of information geometry. The dense N-S sensors, while numerous, are highly correlated in terms of the constraints they provide, and adding more yields diminishing returns. SourceNet has identified this information bottleneck (Tishby et al., 2000). Its attention mechanism acts as a dynamic preconditioner by performing a soft selection that downweights redundant observations to mitigate multicollinearity, while prioritizing the under-sampled orthogonal views (E-W) to stabilize the inversion. This demonstrates that the model has implicitly found the D-Optimality design (Pukelsheim, 2006; Krause et al., 2008), i.e. maximizing the determinant of the Fisher Information Matrix by focusing on the most variance-reducing observations.

### v2

@@ Abstract
Inferring high-dimensional physical states from sparse, ad-hoc sensor arrays is a fundamental challenge across AI for Science, yet standard architectures like CNNs and DeepSets struggle to capture the irregular geometries and relational physics inherent to domains like seismology. To address this, we propose SourceNet, a Transformer-based framework that bridges the profound Sim-to-Real gap via Physics-Structured Domain Randomization (PSDR), a protocol that randomizes governing physical dynamics to enforce invariance to unmodeled environmental heterogeneity. By pre-training on 100,000 synthetic events and fine-tuning on $\sim$2,500 real-world events, SourceNet achieves state-of-the-art precision on held-out real data, demonstrating exceptional data efficiency and real-time capability compared to classical solvers. Beyond prediction, interpretability analysis reveals that the model shows scientific-agent-like features: it autonomously discovers geometric information bottlenecks and learns an attention policy that prioritizes sparse sensor placements, effectively recovering principles of optimal experimental design from data alone.
@@ 1 Introduction
How do we reconstruct a complex physical event, whether it is a distant earthquake, a cosmic radio burst, or a climate anomaly, using only a handful of scattered sensors? [^1] Code and data are available at: https://github.com/jiazhe868/SourceNet This is the archetypal challenge across the physical sciences (Bergen et al., 2019; Bouman et al., 2016; Ravuri et al., 2021): inferring complex physical source parameters $\theta$ from a sparse set of indirect, noisy observations $X=\{x_{1},\dots,x_{n}\}$. Unlike images or video where data pixels are fixed on a regular grid, scientific data is often collected by ad-hoc sensor arrays where the number of sensors varies per event and their spatial arrangement is irregular.
@@
Compounding the geometric challenge is the Sim-to-Real gap (Tobin et al., 2017). In science disciplines like geophysics and robotics, obtaining ground-truth labels for real-world observations is expensive and labor-intensive, therefore models have to be trained on physics-based simulations, which inevitably lack the dirty reality including instrumental noise, realistic wave propagation effect, and structural heterogeneity (Karniadakis et al., 2021). Hence, naive training on idealized simulations faces a substantial distribution shift and is difficult to generalize to various real-world data. A trustworthy AI for Science framework must learn invariant physical operations that hold true despite these distribution shifts.
@@
    We bridge the Sim-to-Real gap not by feature alignment, but by randomizing the physics of the generating process (e.g., velocity models, sensor dropouts), forcing the model to learn representations invariant to crustal heterogeneity. We validate that the gap is bridged through manifold alignment.
@@ 2.1 Deep Learning on Sets and Irregular Grids
Handling varying cardinality inputs requires permutation-invariant architectures. Early set-based architectures, such as DeepSets (Zaheer et al., 2017) and PointNet (Qi et al., 2017), can achieve permutation invariance by applying independent encoders followed by symmetric global pooling. While effective for simple classification tasks, we argue the global pooling operation is too lossy for wave physics, because it reduces the set information using set $\mathrm{sum}$ or $\mathrm{max}$, which limits the network’s ability to model relational context (e.g., amplitude differences between stations) critical for inversion. Graph Neural Networks (GNNs) (Kipf, 2016; Battaglia et al., 2018) model dependencies explicitly but typically assume fixed edges or require expensive dynamic graph construction ($O(N^{2})$) (Wang et al., 2019). More recently, Neural Operators (e.g., DeepONet(Lu et al., 2021), FNO(Li et al., 2020)) have emerged to learn mappings between infinite-dimensional function spaces. However, these methods often assume dense sampling of fixed control points. In contrast, Set Transformers (Lee et al., 2019) leverage self-attention to model pairwise interactions dynamically. In the context of wave physics, attention-based mechanisms offer a theoretical advantage over pooling-based baselines by preserving radiation patterns on sparse arrays.
@@ 2.2 Sim-to-Real Transfer and Domain Randomization
Bridging the simulation-to-reality gap is a key challenge of domains ranging from robotics (Tobin et al., 2017) and neurobioscience (Gonçalves et al., 2020) to geophysics. In these fields, simulations provide abundant data but inevitably simplify the complex, unmodeled dynamics of the physical world (Pan & Yang, 2009). Traditional unsupervised domain adaptation methods aim to align feature distributions via adversarial training (Ganin et al., 2016) or minimizing maximum mean discrepancy (Long et al., 2015), but these require direct access to the target domain during training and can be unstable. Domain randomization (DR) (Tobin et al., 2017) takes a different approach: it randomizes simulator parameters to force the model to learn invariants (Peng et al., 2018). While DR has become standard in robotics for handling visual and contact dynamics (Akkaya et al., 2019), scientific inverse problems require a deeper form of adaptation because the uncertainty lies in the physics themselves. Our SourceNet extends DR by randomizing the governing physical processes (velocity structure, wave propagation effect, scattering) rather than just visual appearance. This not only solves the specific challenge of seismic inversion but offers a generalizable blueprint for learning robust operators in many physical systems with imperfect simulations.
@@ 3 Problem Formulation and Methodology
The physical source ($\mathbf{y}$). Our target is the seismic event $\mathcal{E}$ parameterized by a latent physical state $\mathbf{y}\in\mathbb{R}^{6}$. This state encapsulates the Moment Tensor ($3\times 3$ symmetric tensor) (Aki & Richards, 2002), representing the fault orientation and slip direction (5 independent zero-trace $M_{ij}$ components), and the moment magnitude $M_{W}$ indicating the energy scale (Kanamori, 1977).
The Observation Set ($\mathbf{X}$). Unlike images defined on regular grids, a seismic network is an ad-hoc collection of sensors. We represent the observation as an unordered set $\mathbf{X}=\{(\mathbf{w}_{i},\mathbf{s}_{i})\}_{i=1}^{N}$, where $N$ is the number of stations, which varies per event. One of the key motivations for requiring the inverse mapping to be permutation-invariant is this varying nature of $N$, which will mess up most fixed representations of the data. Each element in the set is a multi-modal tuple: $\mathbf{w}_{i}\in\mathbb{R}^{C\times T}$ represents the dual-domain waveform features, including both the raw velocity time-series (for phase information) and their spectral representations (for frequency-dependent properties); $\mathbf{s}_{i}\in\mathbb{R}^{d}$ represents scalar metadata including station azimuth, epicentral distance, locations, and the amplitude and amplitude ratios of the waveforms. Thus, they encode the physical attenuation from distance-based amplitude decay necessary to unravel the source magnitude.
@@ 3.1 Physics-Structured Domain Randomization (PSDR)
  • Network Availability ($\mathcal{M}$): Real sensors can fail or have gaps in data collection. We apply a stochastic masking operator $\mathcal{M}$ (Bernoulli dropout) to randomly discard stations during training, to make the model robust to sparse and variable geometries.
Validation via Manifold Alignment. To validate our physics-structured approach, we compare the latent spaces of the baseline model trained on clean synthetics and SourceNet trained on PSDR, where both models are fine-tuned on the same subset of real data. As shown in the t-SNE (Maaten & Hinton, 2008) visualizations in Figure 3, standard fine-tuning is insufficient to bridge the reality gap, shown by the disjoint clusters of the naive baseline model, and it memorizes real samples rather than learning generalized physics. In contrast, SourceNet achieves perfect manifold alignment. This proves that PSDR is crucial for effective transfer learning, allowing the model to learn domain-invariant physical features that fine-tuning alone cannot recover.
@@ 3.2 SourceNet Architecture
SourceNet is designed to mirror how a seismologist would analyze an event: first extracting features from each station (phases, amplitudes, location), and then synthesizing this information globally to constrain the source. As illustrated in Figure 1, the model consists of three functional modules:
@@
  • P-Wave Tower: We extract a 6-second window ([1s,+5s][-1s,+5s] relative to the P wave onset) filtered in the low-frequency band (0.1–2.0 Hz) to remove the drift and noise. This stream processes both the raw velocity waveforms and their spectral representations via a 1D ResNet (He et al., 2016). This stream captures not only the high-frequency onset polarity, but also the early crustal phases in the coda. As our analysis reveals, these converted phases contain rich constraints often ignored by traditional first-motion solvers.
@@
  • Scalar Tower: Standard Transformers (Vaswani et al., 2017) use discrete positional encodings (takes input token index within a sentence), which fails for variable and irregular sensor arrays. We instead project a vector of 20 explicit physical features through an MLP. This vector includes the geometric metadata: Station coordinates (lat/lon), azimuth, epicentral distance, and source depth; the amplitude information: max absolute amplitudes for both time-domain and spectral channels (6 for P-wave, 6 for S-wave); and 3 P/S amplitude ratios. This generates a continuous embedding that allows the scale-invariant Transformer to recover the absolute Moment Magnitude ($M_{W}$) and the source radiation patterns.
@@
2. Global Context via Self-Attention. A challenge to the source inversion is that local observations are ambiguous: a single sensor cannot distinguish between different physical parameters without reference to other sensor data. Standard pooling (DeepSets) can fail because it processes these ambiguous features independently and identically for each sensor and then uses a global pooling operation, which masks the relational context.
We employ a Transformer Encoder to model these relations. Through self-attention, each station queries the entire array, aggregating information from distant sensors to update its own representation. This process contextualizes the local data features against the global radiation pattern, which allows each station’s embedding to encode not just what it saw, but how it fits into the collective source mechanism. The simplest example of a helpful pairwise feature is taking a distance estimate using the lat/lon coordinates for two sensors.
@@ 4.1 Datasets
Real-world Fine-tuning Domain. We use a high-quality catalog of 2,435 $M>3.0$ earthquakes in Southern California with focal mechanism solutions from (Yang et al., 2012). These events are selected for their reliability (Quality A & B) and acceptable station coverage (>>5 stations) (Hardebeck & Shearer, 2002).
@@ 5 Experimental Results and Analysis
We evaluate SourceNet on both the held-out synthetic test set (for theoretical upper bound) and the real-world Southern California catalog (to measure Sim-to-Real transfer quality).
@@ 5.1 Sim-to-Real Generalization
As shown in Appendix B, SourceNet predicts source parameters very well on the synthetic test set, with Mean Absolute Errors (MAE) for the moment tensor ($M_{ij}$) ranging from $0.05$ to $0.08$. This shows that the architecture can well approximate the nonlinear wave physics. When applied to real-world data (Figure 4, Bottom), we observe a degradation due to the inevitable reality gap. However, the predictions remain nicely correlated with ground truth, and the magnitude prediction remains robust ($MAE=0.11$) even though real-world waveforms are distorted by scattering. As we demonstrate in Section 5.3, this robustness is directly attributable to our multi-modal design, which disentangles geometric spreading (via the scalar tower) from source physics.
Figure 4: Sim-to-Real Inference Performance. Evaluation on the held-out real-world catalog. (a) Predicted vs. true Magnitude ($M_{W}$) demonstrates high accuracy ($MAE=0.11$), confirming that the scalar tower successfully disentangles source energy from path attenuation. (b-c) Normalized Moment Tensor components ($M_{xx},M_{xy}$) maintain robust linearity despite the domain gap. (d) The Kagan angle error distribution (Mean $23.9^{\circ}$, Median $19.7^{\circ}$) confirms that SourceNet achieves SOTA precision and effectively hits the label uncertainty floor of the manual catalogs.
@@ 5.2 Benchmarking against SOTA
We benchmark SourceNet against established baselines: the authoritative Southern California catalog (Yang et al., 2012) and recent deep learning solvers. On our operational test set that mixes both good (quality A) and acceptable (quality B) events, SourceNet achieves a mean Kagan angle of $23.9^{\circ}$ and a median of $19.7^{\circ}$. Given that ground-truth uncertainties for Quality B events range up to $35^{\circ}$ (Yang et al., 2012), our error effectively hits the label noise floor. This suggests the model is not merely fitting noisy labels but has learned a robust physical operator more consistent than the noisy manual labels used for supervision.
@@ 5.3 Ablation Studies
To validate our architecture, we perform component-wise ablation studies on the synthetic test set. We train two variants: one removing the Scalar Tower (waveform only), and one replacing the Set Transformer with a DeepSets-style pooling architecture. The results are summarized in Table 1. We found that removing the scalar metadata causes the Magnitude prediction to collapse as MAE increasing by over 150% (0.07 $\to$ 0.18). This shows that waveform shapes alone are ambiguous on energy scale, and explicit geometric context (distance, attenuation) fed by the Scalar Tower is essential for disentangling energy from propagation effects. Also, we found replacing the Transformer with simple pooling (DeepSets) makes the mechanism estimation worse: the $M_{ij}$ error nearly triples (0.06 $\to$ 0.17), and the Kagan angle degrades significantly ($6.6^{\circ}\to 28.4^{\circ}$). This suggests that wave physics is relational: determining a source mechanism requires comparing relative data information across the array. Self-Attention captures these pairwise dependencies, whereas global pooling masks them.
@@ 6 Interpretability: Emergent Physical Strategies
Deep learning models are often criticized as black boxes (Rudin, 2019). Here, we show that SourceNet exhibits useful properties beyond simply fitting a function, and has autonomously discovered physical strategies that match human intuitions (Iten et al., 2020).
A question of interest is how much of a waveform is actually useful. Classical solvers typically rely on a single data point: the first P-wave polarity. The rest of the time series is often discarded as coda noise due to complex scattering.
However, our Grad-CAM analysis (Figure 5) shows that SourceNet’s attention is not confined to the P-wave onset but extends substantially into the coda and early scattered phases. Physically, this suggests the model has learned that these phases are not random but contains rich information about the source mechanism (Aki & Richards, 2002). Thus, SourceNet effectively performs an implicit full-waveform inversion from the entire signal duration that human analysts usually ignore (Fichtner, 2010).
@@
This behavior demonstrates an autonomous discovery of information geometry. The dense N-S sensors, while numerous, are highly correlated in terms of the constraints they provide, and adding more yields diminishing returns. SourceNet has identified this information bottleneck (Tishby et al., 2000). Its attention mechanism acts as a dynamic preconditioner by performing a soft selection that downweights redundant observations to mitigate multicollinearity, while prioritizing the under-sampled orthogonal views (E-W) to stabilize the inversion. More theoretically explained, the model has implicitly found the D-Optimality design (Pukelsheim, 2006; Krause et al., 2008), i.e. maximizing the determinant of the Fisher Information Matrix by focusing on the most variance-reducing observations.